# The Physics and Metaphysics of Social Powers: Bridging Cognitive Processing and Social Dynamics, a New Perspective on Power through Active Inference

### v1

@@ 4 Conclusion
Hierarchies are inevitable in complex thermodynamic systems. We propose fostering transparent, plastic hierarchies that remain contestable and therefore open to novelty. This approach acknowledges historic and thermodynamic constraints, while acknowledging space for systemic evolution. While hierarchies are inevitable, social scripts should be read-writable by those with an interest in survival within them. This helps reorient our understanding of an unfolding, globalizing, interconnected and interdependent life: participation in a system is only participatory if those involved are able to track the consequences of the variables and thus future state spaces at stake. A true distribution of information processing capabilities is essential for systemic sustainability and proper equitable development. And “if nature is unjust: change nature.” (Xenofeminist manifesto, 2015).

### v2

@@ 4 Conclusion
Hierarchies are inevitable in complex thermodynamic systems. We propose fostering transparent, plastic hierarchies that remain contestable and therefore open to novelty. This approach acknowledges historic and thermodynamic constraints, while acknowledging space for systemic evolution. While hierarchies are inevitable, social scripts should be read-writable by those with an interest in survival within them. This helps reorient our understanding of an unfolding, globalizing, interconnected and interdependent life: participation in a system is only participatory if those involved are able to track the consequences of the variables and thus future state spaces at stake. A true distribution of information processing capabilities is essential for systemic sustainability and proper equitable development. And “if nature is unjust: change nature.” [Laboria Cuboniks, 2015].